OpenStudy (anonymous):

Ann's Coffee Shop makes a blend that is a mixture of two types of coffee. Type A coffee costs Ann $ 5.95 per pound, and type B coffee costs $ 4.10 per pound. This month, Ann made 162 pounds of the blend, for a total cost of $ 786.30 How many pounds of type A coffee did she use?

OpenStudy (anonymous):

؟؟

OpenStudy (anonymous):

We need to set up two equations since we have two unknowns (the weight of A and the weight of B). The first will relate the weight. \[A + B = 162\]The second will relate cost\[5.95 \cdot A + 4.10 \cdot B = 786.30\]